# Privacy-Preserving Instructions for Aligning Large Language Models

### v1

@@ 1 Introduction
Training models to follow instructions is fundamental in the development of large language model (LLM) applications (Wei et al., 2021; Chung et al., 2022; Ouyang et al., 2022). Instructions in the form of prompts provided by users during interactions with LLM applications are often used as training data to refine the backbone models (OpenAI, 2023; Anthropic, 2023). The instructions can be highly sensitive since users may disclose personal information to elicit accurate responses. To illustrate this, we conducted a quantitative analysis on approximately twelve thousand real-world user instructions collected from the Chatbot Arena  (Zheng et al., 2023b). This analysis begins with an LLM-based detection, followed by human verification, as detailed in Appendix B. Our analysis finds at least thousand of sensitive text pieces. A few examples of sensitive instructions are shown in Figure 2.
Given the sensitive nature of user instructions, we identify the following privacy risks in the pipeline of aligning LLMs with instructions, as illustrated in Figure 1. During the data collection stage, potentially sensitive instructions are collected and presented to human annotators (Ouyang et al., 2022). In the subsequent training and deployment stages, trained models may memorize training data, including sensitive instructions, and inadvertently leak this memorized data during deployment. While the privacy risk of model memorization has been discussed in previous studies (Carlini et al., 2021; 2023; Nasr et al., 2023a), to the best of our knowledge, this work is the first to discuss the privacy risks associated with the annotation process.
@@
The standard method for fine-tuning LLMs on private data is to replace standard optimizers with their DP variants (Yu et al., 2021; Li et al., 2022b; Bu et al., 2022; He et al., 2023; Wu et al., 2023). However, merely ensuring that the fine-tuned model satisfies DP can only address the risk of memorization; there is no protection in the data collection stage when the user instructions are exposed to human annotators. To simultaneously address both privacy risks, human annotators and model memorization, we propose using synthetic instructions (generated from privately fine-tuned language models) as substitutes for real user instructions in the alignment pipeline. To support the claim that DP models do not memorize training data (Guo et al., 2022; Stock et al., 2022; Yue et al., 2022), we show text generative models trained without DP may memorize training samples and leak them during inference, while models trained with DP dot not. The results are presented in Appendix C.
Contributions. We make the following contributions.
@@
We introduce a two-stage framework for privately generating high quality synthetic instructions, which can be seamlessly incorporated into the training pipeline of instruction-following models (Figure 3). First, an LLM pre-trained on public data is privately fine-tuned on user-provided instructions, from which a large set of synthetic instructions are sampled (Algorithm 1). We adopt state-of-the-art strategies for DP fine-tuning including initializing with a large pre-trained model and parameter-efficient fine-tuning (Yue et al., 2022; Kurakin et al., 2023).
@@
There is still a large distribution shift between the user-provided instructions and the initial set of private machine-generated instructions. To bridge this gap, we introduce the second stage of a novel private resampling algorithm (Algorithm 2); we select a subset of the initial private synthetic instructions in order to better match the real user instructions. Specifically, we first use a private histogram to approximate the distribution of the real instruction dataset, where each bin corresponds to each cluster learned from synthetic instructions, for example, representing coding questions for a programming language. The construction of the histogram is inspired by MAUVE which uses similar cluster-and-histogram strategy to measure divergences between two datasets Pillutla et al. (2022). The private histogram is then used to filter the initial synthetic instructions.
@@
We demonstrate the performance of the proposed approach on standard methods of aligning language models: for supervised instruction fine-tuning on LLaMA (Touvron et al., 2023a) 7B and 13B models and for reinforcement learning from human feedback on 1.3B Phi-1.5 model (Li et al., 2023b). We evaluate them on instructions from the LMSYS-Chat-1M dataset (Zheng et al., 2023a) and the AlpacaEval benchmark (Dubois et al., 2023). In supervised fine-tuning, training a 7B LLaMA model with resampled DP synthetic instructions leads to an 8.6% relative improvement compared to using the unfiltered initial DP synthetic instructions. This is measured in the win-rate against a baseline model. In reinforcement learning from human feedback, our private approach is comparable to using real instructions without any privacy guarantees.
@@
Overview. We provide background on fine-tuning LLMs on instructions in Section 2 and discuss privacy risks involved. A new framework for generating high-quality synthetic instructions with differential privacy is presented in Section 3. Empirical results demonstrating the utility of those synthetic instructions are provided in Section 4. We end with a conclusion in Section 5.
@@ 2 Privacy Risks and Background
Figure 2: Samples of sensitive instructions in the Chatbot Arena Conversations dataset (Zheng et al., 2023b). We mask the sensitive texts to protect user privacy.
@@
Only recently has it become possible to identify this vulnerability, following the release of several datasets featuring real-world user interactions with instruction-following models (Zheng et al., 2023a; b; Zhao et al., 2024). We focus on two such datasets collected from the Chatbot Arena (Zheng et al., 2023a; b), carefully adhering to their terms of use, and study privacy implications of using real-world instructions.
@@
Private deep learning. We use ($\varepsilon$$\delta$)-differential privacy (DP) (Dwork et al., 2006a) defined as,
@@
Neighboring datasets are defined by adding/removing one sample, where each sample constitutes a single user instruction (text prompt), i.e., example-level differential privacy (Ponomareva et al., 2023, Section 5.1). $\mathcal{M}$ is a (randomized) training algorithm that outputs trained model weights. The most common approach for private deep learning is to clip and add noise to each gradient update (Song et al., 2013; Abadi et al., 2016; Kairouz et al., 2021). We use a variant of DP-SGD (Abadi et al., 2016) with Adam optimizer, which is called DP-Adam (Li et al., 2022b).
Differentially private synthetic data. A widely adopted method for generating private synthetic text in practice is to train a language model on private data with DP, and then repeatedly sample the DP model to generate synthetic text sequences (Augenstein et al., 2019; Bommasani et al., 2019; Putta et al., 2022; Mattern et al., 2022; Yue et al., 2022; Kurakin et al., 2023). In this work, we follow this general approach to generate initial synthetic instructions. Importantly, we extend beyond the scope of prior studies, which primarily focus on simple tasks such as text classification, by investigating the use of private synthetic text in the training of instruction-following models. Moreover, we introduce a novel filtering algorithm that substantially reduce the distributional gap between real and private synthetic texts.
@@ 3 Generating Synthetic Instructions with Differential Privacy
To mitigate the privacy concerns in aligning LLMs with user instructions, we propose a novel privacy-preserving approach by generating synthetic instructions with differential privacy (see Figure 3 for an overview). First, an LLM generator is fine-tuned with DP on user-provided instructions, which is subsequently used to generate a large set of initial synthetic instructions (Algorithm 1). Next, these initial synthetic instructions are resampled according to a private histogram to approximately match the distribution of the real instructions (Algorithm 2). We carefully allocate the privacy budget in Algorithms 1 and 2, and the formal DP guarantees of the entire process can be composed (Kairouz et al., 2015). The generated DP synthetic data can be used for aligning LLMs and the same DP guarantees are applied because of the post-processing property of DP (Dwork et al., 2014).
@@ 3.1 Stage 1: DP Instruction Generator
We take an LLM, pre-trained on public data, and fine-tune on private instruction data using DP-Adam. This is used to generate samples that comprise a large set of initial instruction dataset, see Algorithm 1. Concretely, LLaMA models (Touvron et al., 2023a) are privately fine-tuned on user instructions with LoRA (Hu et al., 2022). LoRA has been shown to have superior privacy-utility trade-offs in fine-tuning of DP text generators (Kurakin et al., 2023).
@@
Since the fine-tuned model itself is private, we can sample as many instructions as we want. However, there is a non-negligible distributional gap between the initial synthetic instructions and real instructions, which increases as the DP guarantee strengthens, likely due to the noise added in DP training, e.g., Table 1 in Section 4.2. Next, the quality of the synthetic instructions can be improved by resampling.
@@ 3.2 Stage 2: Resample with DP Histogram
Our approach is motivated by how synthetic datasets that have higher MAUVE scores Pillutla et al. (2022) with respect to the real dataset achieves better utility (Kurakin et al., 2023). MAUVE is based on first clustering the samples in a dataset into a histogram, and computing the divergence between the two histograms of the two datasets.
Similarly, Algorithm 2 first groups synthetic instructions, ${\mathbb{D}}_{syn}$, into $K$ clusters, $\{{\bm{c}}_{j}\}_{j=1}^{K}$, in the embedding space. Clustering synthetic instructions does not incur additional privacy cost because it constitutes a post-processing step of Algorithm 1 (Dwork et al., 2014). The cluster centroids of ${\mathbb{D}}_{syn}$ are used to create a histogram of real instruction samples in ${\mathbb{D}}_{real}$ where each bin corresponds to each cluster. This histogram, $\bm{h}$}, is made private by adding appropriate Gaussian noise to each bin. The synthetic data, ${\mathbb{D}}_{syn}$, is resampled to match this private histogram. Table 1 in Section 4.2 demonstrates that the filtered ${\mathbb{D}}_{syn}^{\prime}$ is closer to ${\mathbb{D}}_{real}$ than the initial ${\mathbb{D}}_{syn}$, which in turn improves the performance of the model fine-tuned on the filtered dataset. Further studying the clusters of ${\mathbb{D}}_{syn}$ reveals that, for example, the top voted cluster contains instructions of coding questions in Python as shown in Figure 17 of Appendix G. On the other hand, clusters receive least votes do not exhibit any coherent semantics.
@@
The number of clusters, $K$, traverses a trade-off between signal and noise in the DP filtering: a large $K$ gives a fine-grained characterization of the real instruction distribution, but also makes the histogram harder to privatize because of the increased dimension and reduced votes in each cluster. As the votes from real instructions sum up to $N$, a large $K$ reduces the number of votes per-cluster. Therefore, adding proper noise to the histogram for meaningful DP guarantees would result in an inaccurate estimation of the real instruction distribution due to poor signal-to-noise ratio. One good practice for determining the value of $K$ is to ensure that the average number of votes per cluster is sufficiently robust to DP noise. In our experiments, we investigate the quality of resampled synthetic data, measured by MAUVE (Pillutla et al., 2022), for varying $K$. As depicted in Figure 5 of Section 4.2, while the resampling process enhances the quality of synthetic data for all choices of $K$, the improvement initially increases with $K$ but beyond a certain point, it begins to either plateau or decline.
@@ 4 Experiments
We conduct extensive experiments to evaluate our algorithms. In Section 4.2, we assess the distributional gap between synthetic and real instructions. We then investigate the utility of synthetic instructions in two downstream tasks. Specifically, in Section 4.3, we experiment with supervised instruction fine-tuning. In Section 4.4, we experiment with reinforcement learning from human feedback.
@@ 4.1 Setup for Generating Synthetic Instructions
Instruction datasets. We use the LMSYS-Chat-1M dataset as the private dataset (Zheng et al., 2023a), which has one million conversations between real users and instruction-following models collected from April to August 2023. We only use the user instructions and disregard the corresponding machine-generated responses. This data is first preprocessed by deduplication and filtering out non-English instructions, among other steps, as detailed in Appendix E. After preprocessing, approximately 210,000 instructions remain, which we refer to as Chatbot Arena instructions. The Chatbot Arena instructions are then divided into three subsets: 180,000 for the training set, 10,000 for the validation set, and 20,000 for the test set.
We compare the models fine-tuned on private instructions (with DP) with models fine-tuned on out-of-distribution public instructions (without DP). For this, we take instructions from the FLAN dataset (Chung et al., 2022). FLAN comprises question/answer pairs derived from a rich collection of classic NLP tasks. For our purposes, we retain only the instructions. We randomly sample 180,000 instructions from FLAN unless otherwise specified.
@@
Setup for Algorithm 1. The DP generators are fine-tuned from LLaMA models (Touvron et al., 2023a). This ensures that no private data is leaked in the pre-training, since the cutoff date for the pre-training data of LLaMA is prior to the start of the collection of Chatbot Arena instructions. We tune the hyperparameters based on MAUVE scores and undertake several measures to generate high-quality initial synthetic instructions. Details regarding the hyperparameters for both fine-tuning and sampling are in Appendix F.2.
Using MAUVE score between Chatbot Arena instructions and the initial synthetic instructions, we study the effect of model size with LLaMA 7B and 13B models. The results suggest that a larger model synthesizes better quality data (Appendix D.2). One unexpected by-product of gradient clipping in DP fine-tuning is that the clipping threshold has a significant impact on the length of synthetic instructions and, consequently, affects the MAUVE scores. The results of using varying clipping thresholds are in Appendix D.1. In the rest of this paper, generator models are fine-tuned from LLaMA 13B unless otherwise specified. We sample one million initial synthetic instructions from each fine-tune model.

### v2

@@ Abstract
Proceedings of the $\mathit{41}^{st}$ International Conference on Machine Learning, Vienna, Austria. PMLR 235, 2024. Copyright 2024 by the author(s).
@@ 1 Introduction
Training models to follow instructions is fundamental to the development of large language model (LLM) applications (Wei et al., 2021; Chung et al., 2022; Ouyang et al., 2022). Instructions in the form of prompts provided by users during interactions with LLM applications are used to refine the backbone models (OpenAI, 2023). These instructions are often reviewed and annotated by human annotators to create the training data for this refinement process (Tan & Cabato, 2023). However, user-provided instructions can be highly sensitive since users may disclose personal information to elicit accurate responses. To illustrate this, we conducted a quantitative analysis on approximately twelve thousand real-world user instructions collected from the Chatbot Arena  (Zheng et al., 2023b). This analysis begins with an LLM-based detection, followed by human verification, as detailed in Appendix B. Our analysis finds at least thousand of sensitive text pieces. A few examples of sensitive instructions are shown in Figure 2.
Given the sensitive nature of user instructions, we identify the following privacy risks in the pipeline of aligning LLMs with instructions, as illustrated in Figure 1. During the data collection stage, potentially sensitive instructions are collected and presented to human annotators. In the subsequent training and deployment stages, trained models may memorize training data, including sensitive instructions, and inadvertently leak this memorized data during deployment. While the privacy risk of model memorization has been discussed in previous studies (Carlini et al., 2021, 2023; Nasr et al., 2023a; Wu et al., 2024a), to the best of our knowledge, this work is the first to discuss the privacy risks associated with the annotation process.
@@
The standard method for fine-tuning LLMs on private data is to replace standard optimizers with their DP variants (Yu et al., 2021; Li et al., 2022b; Bu et al., 2022; Wu et al., 2024a). However, merely ensuring that the fine-tuned model satisfies DP can only address the risk of memorization; there is no protection in the data collection stage when the user instructions are exposed to annotators. To simultaneously address both privacy risks, human annotators and model memorization, we propose using synthetic instructions (generated from privately fine-tuned language models) as substitutes for real user instructions in the alignment pipeline.
Contributions. Our main contributions are summarized as follows [^1] Our code is available at https://github.com/google-research/google-research/tree/master/dp_instructions..
@@
We introduce a two-stage framework for privately generating high quality synthetic instructions, which can be seamlessly incorporated into the training pipeline of instruction-following models (Figure 3). First, an LLM pre-trained on public data is privately fine-tuned on user-provided instructions, from which a large set of synthetic instructions are sampled (Algorithm 1). We adopt state-of-the-art strategies for DP fine-tuning including initializing with a large pre-trained model and parameter-efficient fine-tuning (Yu et al., 2022; Yue et al., 2022; Kurakin et al., 2023).
@@
There is still a large distribution shift between the user-provided instructions and the initial set of private machine-generated instructions. To bridge this gap, we introduce the second stage of a novel private resampling algorithm (Algorithm 2); we select a subset of the initial private synthetic instructions in order to better match the real user instructions. Specifically, we first use a private histogram to approximate the distribution of the real instruction dataset, where each bin corresponds to each cluster learned from synthetic instructions, for example, representing coding questions for a programming language. The private histogram is then used to filter the initial synthetic instructions. The construction of the histogram is inspired by Pillutla et al. (2022) which uses similar cluster-and-histogram strategy to measure divergences between two datasets.
@@
We demonstrate the performance of the proposed approach on standard methods of aligning language models: for supervised instruction fine-tuning on LLaMA (Touvron et al., 2023a) 7B and 13B models and for reinforcement learning from human feedback on 1.3B Phi-1.5 model (Li et al., 2023b). We evaluate them on instructions from the LMSYS-Chat-1M dataset (Zheng et al., 2023a) and the AlpacaEval benchmark (Dubois et al., 2023). In supervised fine-tuning, training a 7B LLaMA model with resampled DP synthetic instructions leads to an 8.6% relative improvement compared to using the unfiltered initial DP synthetic instructions. The improvement is measured in the win-rate against a baseline model. In reinforcement learning from human feedback, our private approach is comparable to using real instructions without any privacy guarantees.
@@
Overview. We provide background on fine-tuning LLMs on instructions in Section 2 and discuss privacy risks involved. A new framework for generating high-quality synthetic instructions with DP is presented in Section 3. Empirical results demonstrating the utility of those synthetic instructions are provided in Section 4. We end with a conclusion in Section 5.
@@ 2 Privacy Risks and Background
Figure 2: Samples of instructions containing personal information. We mask the sensitive texts to protect user privacy.
@@
Only recently has it become possible to identify this vulnerability, following the release of several datasets featuring real-world user interactions with instruction-following models (Zheng et al., 2023a; Zhao et al., 2024). We focus on two such datasets collected from the Chatbot Arena (Zheng et al., 2023a, b), carefully adhering to their terms of use, and study privacy implications of using real-world instructions.
@@
Private deep learning. We use ($\varepsilon$$\delta$)-DP (Dwork et al., 2006a) defined as,
@@
Neighboring datasets are defined by adding/removing one sample, where each sample constitutes a single user instruction (text prompt), i.e., example-level DP (Ponomareva et al., 2023, Section 5.1). $\mathcal{M}$ is a (randomized) training algorithm that outputs trained model weights. The most common approach for private deep learning is to clip and add noise to each gradient update (Song et al., 2013; Abadi et al., 2016; Kairouz et al., 2021). We use a variant of DP-SGD (Abadi et al., 2016) with Adam optimizer, which is called DP-Adam (Li et al., 2022b).
Differentially private synthetic data. A widely adopted method for generating private synthetic text in practice is to train a language model on private data with DP, and then repeatedly sample the DP model to generate synthetic text sequences (Augenstein et al., 2019; Bommasani et al., 2019; Putta et al., 2022; Mattern et al., 2022; Yue et al., 2022; Kurakin et al., 2023). In this work, we follow this general approach to generate initial synthetic instructions. Importantly, we extend beyond the scope of prior studies, which primarily focus on simple tasks such as text classification, by investigating the use of private synthetic text in the training of instruction-following models. Moreover, we introduce a novel filtering algorithm that substantially reduces the distributional gap between real and private synthetic texts.
Privacy-preserving In-Context Learning (ICL). In addition to aligning LLMs through fine-tuning, ICL is another paradigm for adapting LLMs to specific domains. ICL improves a model’s performance in a target domain by providing instructions and few-shot demonstrations directly in the prompt (Brown et al., 2020). To protect the privacy of in-context prompts (Duan et al., 2023b), a recent line of research focuses on generating differentially private prompts for in-context learning (Duan et al., 2024; Wu et al., 2024b; Tang et al., 2024b; Hong et al., 2024). Our work differs from DP-ICL in two key aspects. Firstly, regarding the threat model, DP-ICL protects sensitive information within in-context prompts, while our approach protects the privacy of training data against both model memorization and exposure to human annotation. Secondly, in terms of algorithms, our framework produces a large set of synthetic instructions rather than few-shot demonstrations, catering to the requirement of a larger set of data in fine-tuning for LLM alignment.
@@ 3 Generating Synthetic Instructions with Differential Privacy
To mitigate the privacy concerns in aligning LLMs, we propose a novel privacy-preserving approach by generating synthetic instructions with DP (see Figure 3 for an overview). First, an LLM generator is fine-tuned with DP on user-provided instructions, which is subsequently used to generate a large set of initial synthetic instructions (Algorithm 1). Next, these initial synthetic instructions are resampled according to a private histogram to approximately match the distribution of the real instructions (Algorithm 2). We carefully allocate the privacy budget in Algorithms 1 and 2, and the formal DP guarantees of the entire process can be composed (Kairouz et al., 2015; Steinke, 2022). The generated DP synthetic data can be used for aligning LLMs and the same DP guarantees are applied because of the post-processing property of DP (Dwork et al., 2014).
@@ 3.1 Stage 1: DP Instruction Generator
We take an LLM, pre-trained on public data, and fine-tune on private instruction data with DP. This is used to generate samples that comprise a large set of initial instruction dataset, see Algorithm 1. Concretely, LLaMA models (Touvron et al., 2023a) are privately fine-tuned on user instructions with LoRA (Hu et al., 2022). LoRA has been shown to have superior privacy-utility trade-offs in fine-tuning of DP text generators (Kurakin et al., 2023). Previous research has shown that DP training is effective at preventing memorization of training data (Guo et al., 2022; Stock et al., 2022; Yue et al., 2022). In Appendix C, we provide further evidence by by showing that text generative models trained without DP can memorize training samples and leak them during inference, whereas DP-trained models do not.
@@
Since the fine-tuned model itself is private, we can sample as many instructions as we want. However, there is a non-negligible distributional gap between the initial synthetic instructions and real instructions, which increases as the DP guarantee strengthens, likely due to the noise added in DP training. The results in Table 1 in Section 4.2 illustrate this phenomenon. We then present a DP resampling algorithm to improve the quality of initial synthetic instructions.
@@ 3.2 Stage 2: Resample with DP Histogram
The pseudocode of our resampling process is presented in Algorithm 2. We first groups synthetic instructions, ${\mathbb{D}}_{syn}$, into $K$ clusters, $\{{\bm{c}}_{j}\}_{j=1}^{K}$, using k-means. Clustering synthetic instructions does not incur additional privacy cost because it constitutes a post-processing step of Algorithm 1. The cluster centroids of ${\mathbb{D}}_{syn}$ are used to create a histogram of real instruction samples in ${\mathbb{D}}_{real}$ where each bin corresponds to each cluster. This histogram, $\bm{h}$}, is made private by adding appropriate Gaussian noise to each bin. The synthetic data, ${\mathbb{D}}_{syn}$, is resampled to match this private histogram. Table 1 in Section 4.2 demonstrates that the filtered ${\mathbb{D}}_{syn}^{\prime}$ is closer to ${\mathbb{D}}_{real}$ than the initial ${\mathbb{D}}_{syn}$, which in turn improves the performance of the model fine-tuned on the filtered dataset. Further studying the clusters of ${\mathbb{D}}_{syn}$ reveals that, for example, the top voted cluster contains instructions of coding questions in Python as shown in Figure 17 of Appendix G. On the other hand, clusters receive least votes do not exhibit any coherent semantics.
Algorithm 2 reduces the gap between real and synthetic datasets by aligning their histograms. The idea of clustering samples in a dataset into a histogram is inspired by MAUVE (Pillutla et al., 2021, 2022). MAUVE is a metric designed to quantify the discrepancy between synthetic and real data. It first selects methods to represent the distributions of both synthetic and real datasets and then measures the divergence frontiers (Djolonga et al., 2020). One specific method involves clustering the samples in a dataset into a histogram and using this histogram as the approximated distribution.
@@
The size of the histogram, $K$, traverses a trade-off between signal and noise: a large $K$ gives a fine-grained characterization of the real instruction distribution, but also makes the histogram harder to privatize because of the increased dimension and reduced votes in each cluster. As the votes from real instructions sum up to $N$, a large $K$ reduces the number of votes per-cluster. Therefore, adding proper noise to the histogram for meaningful DP guarantees would result in an inaccurate estimation of the real instruction distribution due to poor signal-to-noise ratio. One good practice for determining the value of $K$ is to ensure that the average number of votes per cluster is sufficiently robust to DP noise. In our experiments, we investigate the quality of resampled synthetic data, measured by MAUVE, for varying $K$. As depicted in Figure 5 of Section 4.2, while the resampling process enhances the quality of synthetic data for all choices of $K$, the improvement initially increases with $K$ but beyond a certain point, it begins to either plateau or decline.
@@ 4 Experiments
We conduct extensive experiments to evaluate our algorithms. In Section 4.2, we assess the distributional gap between synthetic and real instructions. We then investigate the utility of synthetic instructions in both supervised instruction fine-tuning (Section 4.3) and reinforcement learning from human feedback (Section 4.4).
@@ 4.1 Setup for Generating Synthetic Instructions
Instruction datasets. We use the LMSYS-Chat-1M dataset as the private dataset (Zheng et al., 2023a), which has one million conversations between real users and instruction-following models collected from April to August 2023. We only use the user instructions and disregard the corresponding machine-generated responses. This data is first preprocessed by deduplication and filtering out non-English instructions, among other steps, as detailed in Appendix E. After preprocessing, approximately 200,000 instructions remain, which we refer to as Chatbot Arena instructions. The Chatbot Arena instructions are then divided into three subsets: 180,000 for the training set, 5,000 for the validation set, and the rest for the test set.
We compare the models fine-tuned on private instructions (with DP) with models fine-tuned on out-of-distribution public instructions (without DP). This comparison aims to demonstrate that, despite the abundance of public instruction datasets (Longpre et al., 2023; Lian et al., 2023), in-domain but sensitive user instructions are still essential for aligning LLMs. For this, we take instructions from the FLAN dataset (Chung et al., 2022; Longpre et al., 2023). FLAN comprises question/answer pairs derived from a rich collection of classic NLP tasks. For our purposes, we retain only the instructions. We randomly sample 180,000 instructions from FLAN unless otherwise specified.
@@
Setup for Algorithm 1. The DP generators are fine-tuned from LLaMA models (Touvron et al., 2023a). This ensures that no private data is leaked in the pre-training, since the cutoff date for the pre-training data of LLaMA is prior to the start of the collection of Chatbot Arena instructions. We tune the hyperparameters based on MAUVE scores. Details regarding the hyperparameters for both fine-tuning and sampling are in Appendix F.2.
We undertake several measures to generate high-quality initial synthetic instructions. Using MAUVE score between Chatbot Arena instructions and the initial synthetic instructions, we study the effect of model size with LLaMA 7B and 13B models. The results suggest that a larger model synthesizes better quality data (Appendix D.2). One unexpected by-product of gradient clipping in DP fine-tuning is that the clipping threshold has a significant impact on the length of synthetic instructions and, consequently, affects the MAUVE scores. The results of using varying clipping thresholds are in Appendix D.1. In the rest of this paper, generator models are fine-tuned from LLaMA 13B unless otherwise specified. We sample one million initial synthetic instructions from each fine-tuned model.